# Kazhdan groups of dimension $16$ with an uncountable range of second $\ell^{2}$-Betti numbers

### v1

@@ Abstract
We construct a family of simple, lacunary hyperbolic groups with property $(T)$ that have rational cohomological dimension $16$ and uncountably many values of second $\ell^{2}$-Betti numbers. Along the way, we present new constructions of measurably diverse finitely generated groups, and we prove that the second $\ell^{2}$-Betti number is far from being semi-continuous in the space of marked groups, even assuming good finiteness properties.
@@ 1 Introduction
Let $p\neq q$ be primes. There exists an uncountable family of groups $(G_{\alpha})_{\alpha\in A}$ and an integer $\chi\geq 1$ with the following properties.
@@
    $G_{\alpha}$ is simple, has property $(T)$, and is lacunary hyperbolic.
@@
    $G_{\alpha}$ has cohomological dimension $16$ modulo $\{p,q\}$.
@@
    The second $\ell^{2}$-Betti numbers $b^{(2)}_{2}(G_{\alpha})$ are pairwise distinct. Further, $b^{(2)}_{n}(G_{\alpha})<\infty$ for every $n\in\mathbb{N}$ and every $\alpha\in A$.
@@
    The $\ell^{2}$-Euler characteristic is $\chi^{(2)}(G_{\alpha})=b_{2}^{(2)}(G_{\alpha})+\chi$.
@@
Let $G_{\alpha}$ and $G_{\beta}$ be groups from the family. If they are measure equivalent with index $c>0$, then $b^{(2)}_{n}(G_{\alpha})=cb^{(2)}_{n}(G_{\beta})$ for all $n\in\mathbb{N}$ by Gaboriau’s theorem [Gab02], and thus $\chi^{(2)}(G_{\alpha})=c\chi^{(2)}(G_{\beta})$. Hence
$cb^{(2)}_{2}(G_{\beta})+\chi=b^{(2)}_{2}(G_{a})+\chi=\chi^{(2)}(G_{\alpha})=c\chi^{(2)}(G_{\beta})=cb^{(2)}_{2}(G_{\beta})+c\chi.$
Since $\chi$ is a positive integer, this shows that $c=1$, and thus $b^{(2)}_{2}(G_{\alpha})=b^{(2)}_{2}(G_{\beta})$, so $\alpha=\beta$. ∎
A version of the construction leading to A produces torsion-free hyperbolic groups with prescribed integer second $\ell^{2}$-Betti number (Theorem 5.1). This has the following consequence.
Theorem C.
Let $(n_{i})_{i\in\mathbb{N}}$ be a sequence of natural numbers. There exist property $(T)$ torsion-free hyperbolic groups $G,G_{i}$ of cohomological dimension $16$, with quotient maps $G\to G_{i}$ such that $G_{i}\xrightarrow{i\to\infty}G$ in the space of marked groups, and
$b^{(2)}_{2}(G_{i})=n_{i}\text{ for all }i\in\mathbb{N}\text{ but }b^{(2)}_{2}(G)=0.$
@@
We instead take a different route, and start from a hyperbolic group enjoying the strong rigidity property $(T_{2})$, introduced by Bader and the second author [BS23, BS25]. This group is a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane (Theorem 3.10), which is the source of the number $16$ in our results. Under this condition, building on an observation from [FF25], we will see that the excision theorem from [PS24a] is sufficient to give exact computations of the second $\ell^{2}$-Betti number of iterated group-theoretic Dehn fillings.
@@ Acknowledgements.
FFF is supported by the Herchel Smith Postdoctoral Fellowship Fund. The authors would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Operators, Graphs, Groups, where work on this paper was undertaken. This work was supported by EPSRC grant EP/Z000580/1.
@@ 2 Homology and cohomology
Let $G_{1}\to G_{2}\to\cdots$ be a directed sequence of groups with colimit $G_{\infty}$, and let $M$ be an $R[G_{\infty}]$-module. Then there is an isomorphism
@@
This is the standard fact that homology commutes with filtered colimits. This is of course not true in cohomology, but it does work under some additional assumptions. We point out that the following is only used to compute cohomological dimensions of the groups from A and D. If the reader is happy with the standard estimate of cohomological dimension as one more than the homological dimension [Bie81, Theorem 4.6 on p. 60], then they can use Proposition 2.2, and move to the next section.
@@
Let $G_{1}\to G_{2}\to\cdots$ be a directed sequence of groups with colimit $G_{\infty}$, and let $M$ be an $R[G_{\infty}]$-module. Suppose that the pullback $H^{n}(G_{i+1};M)\to H^{n}(G_{i};M)$ is surjective for all $i\in\mathbb{N}$. Then
@@
Let $X_{1}$ be a model of the classifying space $BG_{1}$111also denoted by $K(G_{1},1)$.. Using mapping cones, we may inductively define $X_{i+1}$ as a model of $BG_{i+1}$ such that the map $X_{i}\to X_{i+1}$ is an inclusion of CW-complexes. The colimit $X$, in this case a union, of the $X_{i}$ is a model of $BG_{\infty}$. The universal covering $\widetilde{X}$, which is a free $G_{\infty}$-CW complex, is a model of $EG_{\infty}$. Let $\bar{X}_{i}$ be the restriction of the universal covering to $X_{i}$. By covering theory we have $\bar{X}_{i}\cong G_{\infty}\times_{G_{i}}\widetilde{X}_{i}$. The cellular cochain complexes
@@
Let $G_{1}\to G_{2}\to\cdots$ be a directed sequence of groups with colimit $G_{\infty}$. Suppose that the following hold for all $i\in\mathbb{N}$.
@@ 3 $\ell^{2}$-Betti numbers of tracial unitary representations
One calls a $\mathcal{N}(\pi)$-linear homomorphism $f\colon M\to N$ of $\mathcal{N}(\pi)$-modules rank dense if for every $x\in N$ there is an increasing sequence of projections $p_{i}\in\mathcal{N}(\pi),i\in\mathbb{N}$, converging to $\mathrm{id}$ such that $p_{i}x\in N$. By [KPV15, Lemma A.5] the embedding $\mathcal{N}(\pi)\hookrightarrow V_{\pi}$ is rank dense because it is dense.
@@
In the proofs of Theorems A and C, we will actually only need that $G$ has property $(T_{2})$: see Remark 5.3.
@@ 4.1 Cohen–Lyndon triples
Let $(G,P,N)$ be a Cohen–Lyndon triple. Then for all $\mathbb{Z}[\bar{G}]$-modules $M$ and all $n\in\mathbb{N}$, the quotient maps $G\to\bar{G}$ and $P\to\bar{P}$ induce isomorphisms
@@
We will only be concerned with an especially well-behaved case.
@@
Let $(G,P,N)$ be a Cohen–Lyndon triple. Suppose that $P$ is virtually cyclic, $N$ is infinite cyclic, and hence $\bar{P}$ is finite. Let $R$ be a ring where $[P:N]$ is invertible. Then for all $R[\bar{G}]$-modules $M$ and all $n\geqslant 3$, the quotient map $G\to\bar{G}$ induces isomorphisms
@@
Let $G$ be hyperbolic relative to $\{P_{\lambda}\}_{\lambda\in\Lambda}$. Let $B\subset G$ be a finite subset. Then for every sufficiently deep normal subgroups $N_{\lambda}<P_{\lambda}$, it holds
@@
[…]
  3. 3.
[…]
We will mostly be concerned with a very particular example of a relatively hyperbolic pair. Let $G$ be a hyperbolic group, and let $g\in G$ be an element of infinite order. Then $g$ is contained in a unique maximal virtually cyclic subgroup denoted $E_{G}(g)$.
@@
Let $G$ be a hyperbolic group, and let $g\in G$ be an element of infinite order. Then $G$ is hyperbolic relative to $E_{G}(g)$.
Putting these results together we obtain the following, which will serve as the inductive step in our constructions.
@@
    The quotient map $G\to\bar{G}$ is injective when restricted to $B$, and induces an embedding $\bar{P}\hookrightarrow\bar{G}$.
@@
By Theorem 4.6, $G$ is hyperbolic relative to $E_{G}(g)$. The Cohen–Lyndon property follows from Theorem 4.4; the hyperbolicity follows from Theorem 4.5, since a group that is hyperbolic relative to a finite subgroup is hyperbolic; the injectivity follows again from Theorem 4.5.
It remains to prove that $\bar{G}$ is non-elementary and without finite normal subgroups. To this end, we find an additional element of infinite order $h\in G$ such that $G$ is hyperbolic relative to $\{E_{G}(g),E_{G}(h)\}$ [Osi07, Lemmas 5.5 and 5.6], see also [Osi06a, Corollaries 1.7 and 4.5]. Consider the normal subgroups $N<E_{G}(g)$ and $\{1\}<E_{G}(h)$, so that the quotient $G\to\bar{G}$ fits into the framework of Theorem 4.5. In particular, $\bar{G}$ is hyperbolic relative to the image of $E_{G}(h)$, which maps injectively and is almost malnormal in $\bar{G}$ by [Osi06b, Theorem 1.4]. If $\bar{G}$ had a non-trivial finite normal subgroup, then a sufficiently high power of $h$ would centralise it, contradicting malnormality. Similarly, if $\bar{G}$ were virtually cyclic, then a sufficiently high power of $h$ would generate a normal subgroup of $\bar{G}$, contradicting malnormality again. ∎
The situation when $E_{G}(g)=\langle g\rangle$ is especially convenient. Following [OT13], we call such elements special.
Proposition 4.8 ([OT13, Proposition 3.4]).
[…]
In the cases above, to obtain desireable properties we needed to pass to a sufficiently deep normal subgroup of the elementary closure of an element. Let us highlight one situation in which this is not necessary.
Lemma 4.9.
@@
    The triple $(G,\langle g\rangle,\langle g\rangle)$ is Cohen–Lyndon;
@@
    The quotient map $G\to\bar{G}$ is injective when restricted to $B$.
@@
This obtained in the course of the proof of [FFS25, Theorem 3.1], the second and third statement are consequences of the theorem, the first one is [FFS25, Lemma 3.5]. ∎

### v2

# Kazhdan groups of dimension $16$ with prescribed second $\ell^{2}$-Betti number

@@ Abstract
We construct a family of simple, lacunary hyperbolic groups with property $(T)$ that have rational cohomological dimension $16$ and whose second $\ell^{2}$-Betti number can be prescribed to be any positive real. Moreover, we construct hyperbolic groups with property $(T)$ whose second $\ell^{2}$-Betti number can be prescribed to be any non-negative rational. Along the way, we present new constructions of measurably diverse finitely generated groups, and we prove that the second $\ell^{2}$-Betti number is far from being semi-continuous in the space of marked groups, even assuming good finiteness properties.
@@ 1 Introduction
There exists an integer $\chi\geqslant 1$ such that the following holds. Let $p\neq q$ be primes. There exists a family of groups $(G_{x})_{x\in\mathbb{R}_{>0}}$ with the following properties.
@@
    $G_{x}$ is simple, has property $(T)$, and is lacunary hyperbolic.
@@
    $G_{x}$ has cohomological dimension $16$ modulo $\{p,q\}$.
@@
    The second $\ell^{2}$-Betti number is $b^{(2)}_{2}(G_{x})=x$. Further, $b^{(2)}_{n}(G_{x})<\infty$ for every $n\in\mathbb{N}$ and every $x\in\mathbb{R}_{>0}$.
@@
    The $\ell^{2}$-Euler characteristic is $\chi^{(2)}(G_{x})=b^{(2)}_{2}(G_{x})+\chi$.
@@
Focusing on rational numbers, a simpler version of the construction leading to A allows to build hyperbolic groups.
Theorem B.
There exists an integer $\chi\geqslant 1$ such that the following holds. There exists a family of groups $(G_{x})_{x\in\mathbb{Q}_{>0}}$ with the following properties.
  1. 1.
    $G_{x}$ has property $(T)$ and is hyperbolic.
  2. 2.
    $G_{x}$ has cohomological dimension $16$ modulo $\pi$, where $\pi$ is the set of prime factors of the denominator of $x\in\mathbb{Q}_{>0}$.
  3. 3.
    The second $\ell^{2}$-Betti number is $b^{(2)}_{2}(G_{x})=x$.
  4. 4.
    The $\ell^{2}$-Euler characteristic is $\chi^{(2)}(G_{x})=b^{(2)}_{2}(G_{x})+\chi$.
In particular, when $x\in\mathbb{N}$, the group $G_{x}$ constructed above has finite cohomological dimension over $\mathbb{Z}$, hence is torsion-free. This construction has the following consequence.
Corollary C.
Let $(m_{i})_{i\in\mathbb{N}}$ be a sequence of natural numbers. There exist property $(T)$ torsion-free hyperbolic groups $G,G_{i}$ of cohomological dimension $16$, with quotients $G\to G_{i}$ such that $G_{i}\xrightarrow{i\to\infty}G$ in the space of marked groups, and
$b^{(2)}_{2}(G_{i})=m_{i}\text{ for all }i\in\mathbb{N}\text{ but }b^{(2)}_{2}(G)=0.$
@@
Let $G_{x}$ and $G_{y}$ be groups from the family. If they are measure equivalent with index $c>0$, then $b^{(2)}_{n}(G_{x})=cb^{(2)}_{n}(G_{y})$ for all $n\in\mathbb{N}$ by Gaboriau’s theorem [Gab02], and thus $\chi^{(2)}(G_{x})=c\chi^{(2)}(G_{y})$. Hence
$cb^{(2)}_{2}(G_{y})+\chi=b^{(2)}_{2}(G_{x})+\chi=\chi^{(2)}(G_{x})=c\chi^{(2)}(G_{y})=cb^{(2)}_{2}(G_{y})+c\chi.$
Since $\chi$ is a positive integer, this shows that $c=1$, and thus $b^{(2)}_{2}(G_{x})=b^{(2)}_{2}(G_{y})$, so $x=y$. ∎
The same argument gives, to the best of our knowledge, the first infinite family of measurably diverse hyperbolic groups with property $(T)$.
Corollary E.
The groups from B are pairwise non-measure equivalent. ∎
@@
We instead take a different route, and start from a hyperbolic group enjoying the strong rigidity property $(T_{2})$, introduced by Bader and the second author [BS23, BS25]. This group is a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane (Theorem 3.10), which is the source of the number $16$ in our results. Under this condition, building on an observation from [FF25], we will see that the excision theorem from [PS24a] is sufficient to give exact computations of the second $\ell^{2}$-Betti number of iterated group-theoretic Dehn fillings. Let us mention that this construction could also be done in the setting of relatively hyperbolic groups, starting from non-cocompact lattices in the octonionic hyperbolic plane, thanks to the recent work [LNP25]; this would produce groups of cohomological dimension $15$ instead.
@@ Acknowledgements.
FFF is supported by the Herchel Smith Postdoctoral Fellowship Fund. The authors thank Antonio López Neumann and Bin Sun for comments on a previous version of this paper. They thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Operators, Graphs, Groups, where work on this paper was undertaken. This work was supported by EPSRC grant EP/Z000580/1.
@@ 2 Homology and cohomology
Let $G_{0}\to G_{1}\to\cdots$ be a directed sequence of groups with colimit $G_{\infty}$, and let $M$ be an $R[G_{\infty}]$-module. Then there is an isomorphism
@@
This is the standard fact that homology commutes with filtered colimits. This is of course not true in cohomology, but it does work under some additional assumptions. We point out that the following is only used to compute cohomological dimensions of the groups from Theorems A and F. If the reader is happy with the standard estimate of cohomological dimension as one more than the homological dimension [Bie81, Theorem 4.6 on p. 60], then they can use Proposition 2.2, and move to the next section.
@@
Let $G_{0}\to G_{1}\to\cdots$ be a directed sequence of groups with colimit $G_{\infty}$, and let $M$ be an $R[G_{\infty}]$-module. Suppose that the pullback $H^{n}(G_{i+1};M)\to H^{n}(G_{i};M)$ is surjective for all $i\in\mathbb{N}$. Then
@@
Let $X_{0}$ be a model of the classifying space $BG_{0}$111also denoted by $K(G_{0},1)$.. Using mapping cones, we may inductively define $X_{i+1}$ as a model of $BG_{i+1}$ such that the map $X_{i}\to X_{i+1}$ is an inclusion of CW-complexes. The colimit $X$, in this case a union, of the $X_{i}$ is a model of $BG_{\infty}$. The universal covering $\widetilde{X}$, which is a free $G_{\infty}$-CW complex, is a model of $EG_{\infty}$. Let $\bar{X}_{i}$ be the restriction of the universal covering to $X_{i}$. By covering theory we have $\bar{X}_{i}\cong G_{\infty}\times_{G_{i}}\widetilde{X}_{i}$. The cellular cochain complexes
@@
Let $G_{0}\to G_{1}\to\cdots$ be a directed sequence of groups with colimit $G_{\infty}$. Suppose that the following hold for all $i\in\mathbb{N}$.
@@ 3 $\ell^{2}$-Betti numbers of tracial unitary representations
One calls a $\mathcal{N}(\pi)$-linear homomorphism $f\colon M\to N$ of $\mathcal{N}(\pi)$-modules rank dense if for every $x\in N$ there is an increasing sequence of projections $p_{i}\in\mathcal{N}(\pi),i\in\mathbb{N}$, converging to $\mathrm{id}$ such that $p_{i}x\in f(M)$. By [KPV15, Lemma A.5] the embedding $\mathcal{N}(\pi)\hookrightarrow V_{\pi}$ is rank dense because it is dense.
@@
Cocompact lattices in the isometry group of the octonionic hyperbolic plane exist [BHC62], and they can be chosen torsion-free up to passing to a finite index subgroup [Sel60]. In the proofs of our main results, we will actually only need that $G$ has property $(T_{2})$: see Remark 5.3.
@@ 4.1 Cohen–Lyndon triples
Let $(G,P,N)$ be a Cohen–Lyndon triple. Then for all $\mathbb{Z}[\bar{G}]$-modules $M$ and all $n\in\mathbb{N}$, the quotients $G\to\bar{G}$ and $P\to\bar{P}$ induce isomorphisms
@@
We will mostly be concerned with an especially well-behaved case.
@@
Let $(G,P,N)$ be a Cohen–Lyndon triple. Suppose that $P$ is virtually cyclic, $N$ is infinite cyclic, and hence $\bar{P}$ is finite. Let $R$ be a ring where $[P:N]$ is invertible. Then for all $R[\bar{G}]$-modules $M$ and all $n\geqslant 3$, the quotient $G\to\bar{G}$ induces isomorphisms
@@
A useful feature of the Cohen–Lyndon property is that it is transitive. The following is a generalisation of [CIOS26, Lemma 4.22], which deals with the case in which $L=N$.
Lemma 4.4.
Let $N<L<P<G$ be groups such that $(G,P,\left<\!\left<N\right>\!\right>_{P})$ and $(P,L,N)$ are Cohen–Lyndon triples (in particular $N$ is normal in $L$). Then $(G,L,N)$ is a Cohen–Lyndon triple.
Proof.
Let $T$ be a set of left transversals of $G/P\left<\!\left<N\right>\!\right>_{G}$ such that $\left<\!\left<N\right>\!\right>_{G}=*_{t\in T}t\left<\!\left<N\right>\!\right>_{P}t^{-1}$, and let $S$ be a set of left transversals of $P/L\left<\!\left<N\right>\!\right>_{P}$ such that $\left<\!\left<N\right>\!\right>_{P}=*_{s\in S}sNs^{-1}$. Hence
$\left<\!\left<N\right>\!\right>_{G}=*_{t\in T}t\left<\!\left<N\right>\!\right>_{P}t^{-1}=*_{t\in T,s\in S}tsN(ts)^{-1},$
so it remains to show that $TS$ is a set of left transversals for $G/L\left<\!\left<N\right>\!\right>_{G}$.
For $g\in G$, there exist $t\in T$ and $p\in P$ such that $g\in tp\left<\!\left<N\right>\!\right>_{G}$. Further, there exists $s\in S$ such that $p\in sL\left<\!\left<N\right>\!\right>_{P}$. Hence
$g\in tp\left<\!\left<N\right>\!\right>_{G}\subset t(sL\left<\!\left<N\right>\!\right>_{P})\left<\!\left<N\right>\!\right>_{G}=tsL\left<\!\left<N\right>\!\right>_{G}.$
Suppose that there exist $t_{i}\in T,s_{i}\in S$ such that $t_{1}s_{1}\in t_{2}s_{2}L\left<\!\left<N\right>\!\right>_{G}$. Then
$t_{2}^{-1}t_{1}\in s_{2}L\left<\!\left<N\right>\!\right>_{G}s_{1}^{-1}=s_{2}Ls_{1}^{-1}\left<\!\left<N\right>\!\right>_{G}\subset P\left<\!\left<N\right>\!\right>_{G}.$
Because $T$ is a set of left transversals of $G/P\left<\!\left<N\right>\!\right>_{G}$, we have $t_{1}=t_{2}$, so $s_{2}^{-1}s_{1}\in L\left<\!\left<N\right>\!\right>_{G}$. Let $l\in L$ be such that $l^{-1}s_{2}^{-1}s_{1}\in\left<\!\left<N\right>\!\right>_{G}$. Now $l^{-1}s_{2}^{-1}s_{1}\in P$, and $\left<\!\left<N\right>\!\right>_{G}\cap P=\left<\!\left<N\right>\!\right>_{P}$ because $(G,P,\left<\!\left<N\right>\!\right>_{P})$ is a Cohen–Lyndon triple [Sun20, Proposition 6.1(a)], so $l^{-1}s_{2}^{-1}s_{1}\in\left<\!\left<N\right>\!\right>_{P}$, that is $s_{1}\in s_{2}L\left<\!\left<N\right>\!\right>_{P}$. Because $S$ is a set of left transversals for $P/L\left<\!\left<N\right>\!\right>_{P}$, we have $s_{1}=s_{2}$ and conclude. ∎
@@
[…]
@@
    The quotient $G\to\bar{G}$ is injective on $B$ and induces embedding $\bar{P}\to\bar{G}$.
We will mostly be concerned with a particular example of a relatively hyperbolic pair. Recall that a subgroup $P<G$ is almost malnormal if $gPg^{-1}\cap P$ is finite for all $g\in G\setminus P$.
@@
Let $G$ be a non-elementary hyperbolic group, and let $P<G$ be quasiconvex and almost malnormal. Then $G$ is hyperbolic relative to $P$.
The main example is the following.
Definition 4.8.
Let $G$ be a non-elementary hyperbolic group, and $g\in G$ an element of infinite order. The elementary closure of $g$ is the unique maximal virtually cyclic subgroup $E_{G}(g)<G$ containing $g$. We say that $g\in G$ is special if it has infinite order and $E_{G}(g)=\langle g\rangle$.
Starting from a quasiconvex almost malnormal subgroup, we can build a new one. The following notion is central in much of the literature on small cancellation theory on groups with hyperbolic features [Ols93, Osi10, Hul16].
Definition 4.9.
Let $G$ be a non-elementary hyperbolic group. A subgroup $H<G$ is suitable if it is non-elementary and does not normalise any non-trivial finite subgroup of $G$.
Once again there is a main example we will care about.
Lemma 4.10 ([Osi25, Lemma 2.4]).
Let $G$ be a non-elementary hyperbolic group with no non-trivial finite normal subgroups. Then every non-trivial normal subgroup of $G$ is suitable.
The following, which is the consequence of the more general [AH21, Theorem 6.2], says that we can extend quasiconvex almost malnormal subgroups via elements of suitable subgroups.
Theorem 4.11 ([AH21]).
Let $G$ be a non-elementary hyperbolic group with no non-trivial finite normal subgroups, and let $P<G$ be an infinite-index quasiconvex almost malnormal subgroup. Let $H<G$ be a suitable subgroup. Then there exists an element $h\in H$ such that $\langle P,h\rangle\cong P*\langle h\rangle$ is an infinite-index quasiconvex almost malnormal subgroup.
We can now use these results to prove the two main technical propositions of this section, which will be used as the induction step in our constructions. The first is about Dehn filling with respect to a large power of an arbitrary element.
@@
    The quotient $G\to\bar{G}$ is injective on $B$, and induces an embedding $\bar{P}\hookrightarrow\bar{G}$.
@@
By Theorem 4.7, $G$ is hyperbolic relative to $P$. The Cohen–Lyndon property follows from Theorem 4.5; the hyperbolicity follows from Theorem 4.6, since a group that is hyperbolic relative to a finite subgroup is hyperbolic; the injectivity follows again from Theorem 4.6.
It remains to prove that $\bar{G}$ is non-elementary and without finite normal subgroups. To this end, we use Theorem 4.11 to find an additional element $h\in G$ such that $G$ is hyperbolic relative to $P\coloneqq E_{G}(g)*\langle h\rangle$. We can now see $G/\left<\!\left<N\right>\!\right>_{G}$ as a Dehn filling of $(G,P)$ with respect to the sufficiently deep normal subgroup $\left<\!\left<N\right>\!\right>_{P}<P$. In particular, $\bar{G}$ is hyperbolic relative to $P/\left<\!\left<N\right>\!\right>_{P}\cong\bar{P}*\langle h\rangle$. Since this is not virtually cyclic and has no non-trivial finite normal subgroups, the same is true for $G$, by [DGO17, Theorem 6.14(b)]. ∎
The second result is about Dehn filling with respect to a prescribed power of a well-chosen element.
Proposition 4.13.
@@
    The triple $(G,\langle g\rangle,\langle g^{k}\rangle)$ is Cohen–Lyndon;
@@
    The quotient $G\to\bar{G}$ is injective on $B$, and the image of $g$ has order $k$.
@@
We first find a special element $y\in xH$. By [AMO07, Lemma 3.8], the suitable subgroup $H$ contains a special element $h$. If $x\in H$, then we can set $y=h\in xH$; otherwise [OT13, Lemma 3.2] ensures that $x\langle h\rangle$ contains a special element $y$. So $E_{G}(y)=\langle y\rangle$ is infinite-index, quasiconvex and almost malnormal in $G$. Applying Theorem 4.11 twice, we find elements $z,w\in H$ such that $P\coloneqq\langle y,z,w\rangle$ is a free group of rank $3$ that is infinite-index, quasiconvex and almost malnormal. By Theorem 4.7, $G$ is hyperbolic relative to $P$. We choose
$g\coloneqq yzwz^{2}w^{2}\cdots z^{n}w^{n},$
where $n$ is large enough such that $g$ is a $C^{\prime}(\lambda)$-small cancellation word in $P$, for some small parameter $\lambda>0$. By classical small cancellation theory (see e.g. [LS01, Chapter V]), the normal subgroup $\left<\!\left<g\right>\!\right>_{P}<P$ is sufficiently deep, hence so is $\left<\!\left<g^{k}\right>\!\right>_{P}$. We can thus apply Theorem 4.6 and obtain that $G/\left<\!\left<g^{k}\right>\!\right>_{G}$ is hyperbolic relative to $P/\left<\!\left<g^{k}\right>\!\right>_{P}$, and that both $B$ and $P/\left<\!\left<g^{k}\right>\!\right>_{P}$ embed into $\bar{G}$.
If $k=1$, then $P/\left<\!\left<g^{k}\right>\!\right>_{P}=P/\left<\!\left<g\right>\!\right>_{P}$ is a classical small cancellation group, hence it is hyperbolic [Gro87]. If $k>1$, then $P/\left<\!\left<g^{k}\right>\!\right>_{P}$ is a one-relator group with torsion, hence Newman’s spelling theorem [New68] implies that it is hyperbolic, and that the image of $g$ has order $k$. In both cases, $\bar{G}$ is hyperbolic relative to a non-elementary hyperbolic group, hence it is itself non-elementary hyperbolic. Since $P/\left<\!\left<g^{k}\right>\!\right>_{P}$ also has no non-trivial finite normal subgroups in both cases, the same is true of $\bar{G}$ by [DGO17, Theorem 6.14(b)]. Finally, in both cases $g$ has order $k$ in $P/\left<\!\left<g^{k}\right>\!\right>_{P}$, which embeds in $\bar{G}$, so $g$ has order $k$ in $\bar{G}$ as well.
It remains to show the first item. Since $\left<\!\left<g^{k}\right>\!\right>_{P}<P$ is sufficiently deep, by Theorem 4.5 the triple $(G,P,\left<\!\left<g^{k}\right>\!\right>_{P})$ is Cohen–Lyndon. Moreover by the Cohen–Lyndon theorem on one-relator groups [CL63], the triple $(P,\langle g\rangle,\langle g^{k}\rangle)$ is Cohen–Lyndon. So Lemma 4.4 implies that $(G,\langle g\rangle,\langle g^{k}\rangle)$ is Cohen–Lyndon. ∎